How accurately can the position of a proton be determined without giving it more than 1.00 keV of kinetic energy?

Grade:12th pass

1 Answers

Arun
6 years ago
v
that the proton has after measurement. Hence KE = p
2
/2mp = 1 keV gives p ≈ 7.31 × 10-
22 N s, which can be taken as ∆px. Hence ∆x = ħ/2(7.31 × 10-22) = 7.22 × 10-14 m.